Equivalent Units of Production and Related Costs The charges to Work in Process-Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in pro production. Work in Process-Assembly Department Bal., 6,000 units, 55% completed Direct materials, 141,000 units $1.3 Direct labor Factory overhead Bal, 2 units, 40% completed 19,350 183,300 To Finished Goods, 138,000 units 378,400 147,140 Determine the following a. The number of units in work in process inventory at the end of the period units . Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in '0. Work in Process-Assembly Department Equivalent Units of Production for Direct Materials and Conversion Costs Whole Units Equivalent Units Direct Materials Equivalent Units Conversion Inventory in process, beginning Started and completed Check My Work 4 more Check My Work uses remaining
b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0". Work in Process-Assembly Department Equivalent Units of Production for Direct Materials and Conversion Costs Whole Units Equivalent Units Direct Materials Equivalent Units Conversion Inventory in process, beginning Started and completed Transferred to finished goods Inventory in process, ending Total units c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent. Costs Per Equivalent Unit Direct Materials Conversion d. Cost of the units started and completed during the period.
Answer #1

a) Ending work in process unit = 6000+141000-138000 = 9000 Units

b) Equivalent unit

Whole Unit EUP-Direct material EUP-Conversion
Beginning work in process 6000 0 6000*45% = 2700
Unit started and completed (138000-6000) 132000 132000 132000
Transferred to finished goods 138000 132000 134700
Ending work in process 9000 9000 9000*40% = 3600
Equivalent unit of production 147000 141000 138300

Cost per equivalent unit :

Direct material = 183300/141000 = 1.3

Conversion = (378400+147140)/138300 = 3.8

Cost of unit started and completed = (1.3+3.8)*132000 = $673200